OpenStudy (anonymous):

Convert the following repeating decimal to a/b form, whrer a, b are integers and b is not = 0 -0.02 there is a line over the 2 but I didn't know how to put that in. Also, this is not a negetive. The line should be raised also.

OpenStudy (cwrw238):

it must be 0.022222222..... that is 2 recurring

OpenStudy (cwrw238):

let x be 0.02.... then 100x = 2.222.... 10x =0.22222 90x = 2.222222 - 0.2222.. = 2 so x = 2/90 = 1/45

OpenStudy (anonymous):

what about -2.34, again not a negative and there should be a line over the 4

OpenStudy (cwrw238):

use similar procedure deal with the decimal part 0.34444.... this equals 0.3 + 0.0444.... = 3/10 + 2/45 = 27/90 + 4/90 = 31/90 so we get 2 31/90

OpenStudy (anonymous):

Thank you so much, this really explained it for me!!